Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability 2

Expert replies
by Gurpinder » Thu Jul 29, 2010 2:09 pm
A bag contains six marbles: two red, two blue, and two green. If two marbles are drawn at random, what is the probability that they are the same color?

oa: [spoiler]1/5[/spoiler]

i keep on getting 1/15???

how should i tackle these pair problems.
Join the discussion
Source: — Problem Solving |

by nithi_mystics » Thu Jul 29, 2010 2:19 pm
Prob. of getting two red balls = 2C2/6C2 = 1/15

p(2 blue balls) = p(2 green balls) = 1/15

Prob of getting 2 balls of the same color = 1/15 + 1/15 + 1/15 = 1/5
Thanks
Nithi
Join the discussion

by JeetGulia » Thu Jul 29, 2010 2:20 pm
Gurpinder wrote:A bag contains six marbles: two red, two blue, and two green. If two marbles are drawn at random, what is the probability that they are the same color?

oa: [spoiler]1/5[/spoiler]

i keep on getting 1/15???

how should i tackle these pair problems.
Total way to select 2 marbles 6!/4!2!=15
Fav way to select 2 marble 3 only 3 set of same color marble.

Prob = 3/15 = 1/5
Join the discussion

by Gurpinder » Thu Jul 29, 2010 2:22 pm
nithi_mystics wrote:Prob. of getting two red balls = 2C2/6C2 = 1/15

p(2 blue balls) = p(2 green balls) = 1/15

Prob of getting 2 balls of the same color = 1/15 + 1/15 + 1/15 = 1/5
but doesent the problem say that you pick 2 balls at random. so doesent that mean that your choice on the first ball will effect the options for the next selection. thats what i was doing.

1/3 for 1st selection.
1/5 for 2nd selection

= 1/15
Join the discussion

by jeffedwards » Thu Jul 29, 2010 3:19 pm
Just think of it this way, it doesn’t matter what color you chose first, so ignore that.
No suppose you chose any color, what is the probability of choosing a match….1/5
Because you have one marble that will match that selection out of a total of 5 marbles left.
Join the discussion

by Gurpinder » Thu Jul 29, 2010 3:30 pm
jeffedwards wrote:Just think of it this way, it doesn’t matter what color you chose first, so ignore that.
No suppose you chose any color, what is the probability of choosing a match….1/5
Because you have one marble that will match that selection out of a total of 5 marbles left.
what if the ratio of color of balls was different?
Join the discussion

by jeffedwards » Thu Jul 29, 2010 4:55 pm
Gurpinder wrote:
jeffedwards wrote:Just think of it this way, it doesn’t matter what color you chose first, so ignore that.
No suppose you chose any color, what is the probability of choosing a match….1/5
Because you have one marble that will match that selection out of a total of 5 marbles left.
what if the ratio of color of balls was different?

Good question

Then you would have to do the probability separate. So in this case you would solve the problem just as nithi_mystics.

Probability of red 2/6 or 1/3
Probability of red again 1/5
Times it (1/3)*(1/5) = 1/15 (That’s the probability of drawing red and red again)
Repeat the process for all three colors and add the results
(1/15)+ (1/15)+ (1/15) = 3/15 = 1/5
1/5 is your answer
Join the discussion